OpenStudy (anonymous):

use logarithmic differentiation to find the derivative of the function.y = x^(9x)

OpenStudy (anonymous):

y = x^9x ln y = ln (x^9x), let's assume x > 0 then y > 0 lny = 9x*ln(x) differentiating leads us to y'/y = 9ln(x) + 9 so y' = x^(9x) * (9ln(x)+9)

OpenStudy (anonymous):

Did you follow that malik? ^^

OpenStudy (anonymous):

kinda...i really starting to hate calc..